Please help with my homework problem #46. Evaluate each polynomial for the given replacements of the variables.
2pq^3 - 5q^2 + 8p, for p = -4 and q = -2
Thank you so much.

Answer by sdmmadam@yahoo.com(530)

2pq^3 - 5q^2 + 8p, for p = -4 and q = -2
Putting p = -4 and q = -2 in
2pq^3 - 5q^2 + 8p
=2X(-4)X(-2)^3 - 5X[(-2)^2] + 8X(-4)
= (-8)X(-8)-5X(4)-32
=64-20-32
= 64 -52
=12
Answer:2pq^3 - 5q^2 + 8p = 12 for p = -4 and q = -2